OpenStudy (anonymous):

What is the approximate perimeter of a semicircle with a radius of 8 m? Use pi = 3.14. Express your final answer to the nearest hundredth.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Data_LG2

rishavraj (rishavraj):

see perimeter of circle = \[2\pi r\] semicircle means half circle u know wht do do next???

OpenStudy (igreen):

\(P = 2\pi r\) Plug in 8 for 'r' and '3.14' for 'pi'. Simplify

OpenStudy (igreen):

Afterwards, we divide by 2, since it's half a circle.

OpenStudy (anonymous):

is 50.24 right?

rishavraj (rishavraj):

no u need to divide it by 2...all bcoz its a semecircle..

OpenStudy (igreen):

Correct, now divide by 2. 50.24 / 2 = ?

OpenStudy (anonymous):

25.12

OpenStudy (igreen):

Correct, that's our answer.

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

Is it ok if i tag you if i need anymore help?

rishavraj (rishavraj):

* nearest hundredth.

OpenStudy (anonymous):

ohh so it would be 25.1?

OpenStudy (igreen):

No..25.12 is correct.

OpenStudy (igreen):

25.12 ^ Tenths place 25.12 ^ Hundredths place

OpenStudy (anonymous):

Ok thanks @iGreen